OpenStudy (anitaasks):

Summation from k=0 to n of [(1/k+4) - 1/(k+5)]?

OpenStudy (michele_laino):

hint: we can write: \[\left( {\frac{1}{4} - \frac{1}{5}} \right) + \left( {\frac{1}{5} - \frac{1}{6}} \right) + \left( {\frac{1}{6} - \frac{1}{7}} \right) + ... + \left( {\frac{1}{{n + 4}} - \frac{1}{{n + 5}}} \right) = ...?\]

OpenStudy (jhannybean):

I see how it works now.

OpenStudy (michele_laino):

hint: so we have: 1/5 - 1/5 = 0, -1/6 + 1/6 = 0 -1/7 + 1/7 = 0 and so on, the only 2 terms remain, namely the first one, and the last one

OpenStudy (michele_laino):

hint: the first term is 1/4, and the last term is: -1/(n+5) so, what is the final result?

OpenStudy (anitaasks):

Ohh I see it now. Thank you!!

OpenStudy (michele_laino):

Thank you!!